### v1

# Evaluating Sparse Autoencoders: From Shallow Design to Matching Pursuit

Abstract
Sparse autoencoders (SAEs) have recently become central tools for interpretability, leveraging dictionary learning principles to extract sparse, interpretable features from neural representations whose underlying structure is typically unknown. This paper evaluates SAEs in a controlled setting using MNIST, which reveals that current shallow architectures implicitly rely on a quasi-orthogonality assumption that limits the ability to extract correlated features. To move beyond this, we introduce a multi-iteration SAE by unrolling Matching Pursuit (MP-SAE), enabling the residual-guided extraction of correlated features that arise in hierarchical settings such as handwritten digit generation while guaranteeing monotonic improvement of the reconstruction as more atoms are selected.
Machine Learning, ICML
1 Introduction
Sparse dictionary learning (Mairal et al., 2009; Rubinstein et al., 2010; Tosic & Frossard, 2011) aims to represent data ${\bm{x}}^{i}$ as sparse linear combinations of learned basis vectors (${\bm{D}}$, a.k.a atoms), for $i=1,\dots,n$, i.e.,
${\bm{x}}^{i}\approx{\bm{D}}{\bm{z}}^{i}\quad\text{subject to}\quad\|{\bm{z}}^{i}\|_{0}\leq k,\quad\forall\ i=1,\dots,n.\vspace{-1.5mm}$ (1)
where the constraint ensures that the sparse code ${\bm{z}}^{i}$ is at most $k$-sparse. Sparse representations are ubiquitous in science and engineering, with applications in medical imaging (Lustig et al., 2007; Hämäläinen et al., 2013), image restorations (Mairal et al., 2007; Dong et al., 2013), transcriptomics (Cleary et al., 2021), and genomics (Lucas et al., 2006) with origins from computational neuroscience (Olshausen & Field, 1997, 1996). The formulation in (1) leads to a bi-level optimization problem: the inner problem performs sparse approximation to estimate the code ${\bm{z}}^{i}$ given the dictionary ${\bm{D}}$, while the outer problem updates the dictionary based on the current code estimates to better represent the data (Tolooshams, 2023).
Solving this bi-level optimization can be achieved via alternating minimization (Agarwal et al., 2016), alternating between the inner and outer problems until a convergence criterion. The inner problem has been extensively studied in the compressed sensing literature (Donoho, 2006; Candès et al., 2006). Classical approaches include greedy $\ell_{0}$-based algorithms (Tropp, 2004) such as Matching Pursuit (Mallat & Zhang, 1993) and Orthogonal Matching Pursuit (Pati et al., 1993), as well as convex relaxation $\ell_{1}$-based methods (Chen et al., 2001; Tibshirani, 1996). Since sparse recovery lacks a closed-form solution (Natarajan, 1995), the sparse approximation step typically requires multiple residual-based iterations to converge.
Prior work solves dictionary learning by optimizing the outer problem using closed-form least squares (Agarwal et al., 2016), local gradient updates (Chatterji & Bartlett, 2017), or sequential methods like MOD (Engan et al., 1999) and K-SVD (Aharon et al., 2006). A key bottleneck lies in the repeated solution of the inner sparse coding problem, which typically converges sublinearly (Beck & Teboulle, 2009; Moreau & Bruna, 2017). To address this, the unrolling literature proposes turning the iterative solver into a neural network, enabling fixed-complexity approximations. This idea, sparked by LISTA (Gregor & LeCun, 2010), has been shown to achieve linear convergence (Chen et al., 2018; Ablin et al., 2019) and accelerate both inference and learning (Tolooshams & Ba, 2022; Malézieux et al., 2022). Unrolling further allows the inner and outer optimization to be directly mapped to forward inference and backward learning in deep networks (Tolooshams et al., 2020).
Sparsity has been established as a useful prior for interpretability (Mairal et al., 2014; Lipton, 2017; Ribeiro et al., 2016). Building on this principle, recent work has proposed the use of sparse autoencoders (SAEs) to extract human-interpretable features from the internal activations of large language models (LLMs) (Elhage et al., 2022; Cunningham et al., 2023; Bricken et al., 2023; Rajamanoharan et al., 2024; Gao et al., 2025). These methods are motivated by the linear representation hypothesis  (Arora et al., 2016; Olah et al., 2020; Park et al., 2024) and the superposition hypothesis  (Elhage et al., 2022), i.e., internal representations can be modeled as sparse linear combinations of semantic directions.
Algorithm 1 Matching Pursuit Sparse Autoencoders (MP-SAE)
  Input: Dictionary ${\bm{D}}$, bias $\bm{b}_{\text{pre}}$, data $\bm{x}$, steps $T$
  Initialize residual: $\bm{r}^{(0)}=\bm{x}-\bm{b}_{\text{pre}}$
  Initialize reconstruction: $\hat{\bm{x}}^{(0)}=\bm{b}_{\text{pre}}$
  Initialize sparse code: $\bm{z}^{(0)}=\bm{0}$
  for $t=1,\ldots,T$ do
     $j^{(t)}=\operatorname*{arg\,max}_{j=\{1,\ldots,p\}}(\bm{{\bm{D}}}^{\top}\bm{r}^{(t-1)})_{j}$
     ${\bm{z}}^{(t)}_{j^{(t)}}={\bm{D}}_{j^{(t)}}^{\top}{\bm{r}}^{(t-1)}$
     $\hat{\bm{x}}^{(t)}=\hat{\bm{x}}^{(t-1)}+{\bm{z}}^{(t)}_{j^{(t)}}{\bm{D}}_{j^{(t)}}$
     ${\bm{r}}^{(t)}={\bm{r}}^{(t-1)}-{\bm{z}}^{(t)}_{j^{(t)}}{\bm{D}}_{j^{(t)}}$
  end for
  Output: Sparse code $\bm{z}=\sum_{t=1}^{T}{\bm{z}}^{(t)}$             Reconstruction $\hat{\bm{x}}=\hat{\bm{x}}^{(T)}$
Figure 1: MP-SAE. Top : Full forward pass. Bottom: One encoder iteration.
Although SAEs are widely used for model interpretability (Kim et al., 2018; Cunningham et al., 2023), they have rarely been evaluated in small-scale, controlled settings, despite their strong ties to classical dictionary learning. Among the few studies, (Hindupur et al., 2025) demonstrate that SAEs impose structural assumptions that shape what they can and cannot detect.
Our Contributions In this work, we revisit the relationship between SAEs and dictionary learning by testing modern architectures on MNIST, a widely used benchmark in sparse coding (Aharon et al., 2006; Makhzani & Frey, 2014). Despite similarities in architecture, we find that SAEs with different sparsity mechanisms yield structurally distinct dictionaries. These differences may have important implications for interpretability.
We demonstrate that shallow SAEs implicitly favor near-orthogonal dictionaries, due to their one-shot sparse inference. To investigate this further, we introduce MP-SAE, which unrolls Matching Pursuit into a sequential, residual-guided inference process that operates effectively in regimes with highly correlated features and is supported by convergence guarantees. MP-SAE learns a globally correlated set of atoms but uses a low-redundancy subset to represent each input, a property missing in current shallow SAEs.
Compared to shallow SAEs, MP-SAE yields more expressive representations and naturally constructs a representational hierarchy—first selecting atoms that capture coarse structure, then adding finer details. This coarse-to-fine behavior may lead to more interpretable representations, as it mirrors the hierarchical organization of real-world features (Bussmann et al., 2025).
2 Background
Representation Hypotheses Efforts to understand neural representations through interpretability are often motivated by two guiding hypotheses: the Linear Representation Hypothesis and the Superposition Hypothesis (Arora et al., 2016; Olah et al., 2020; Park et al., 2024; Elhage et al., 2022). These suggest that internal activations of large deep neural networks can be expressed as linear combinations of human-interpretable concept directions ${\bm{D}}$. In practice, the number of possible concepts $p$ far exceeds the dimensionality $m$ of the representation: $m\ll p$, leading to superposition—multiple concepts overlapping within the same activation (Elhage et al., 2022). Despite this, meaningful disentanglement is possible under the sparsity assumption: $\|\bm{z}\|_{0}\leq k$, where only a small number $k\ll p$ of concepts are active (Donoho, 2006).
Sparse Concept Extraction as Dictionary Learning As formalized in (Fel et al., 2023), the task of concept extraction in interpretability can be cast as a dictionary learning problem : learn a set of interpretable directions ${\bm{D}}$ such that activations $\bm{x}$ can be approximated by sparse linear combinations with sparse code $\bm{z}$ (See equation 1). In practice, this is most often implemented using shallow SAEs, which have been shown to extract meaningful and monosemantic concepts across a variety of architectures and domains (Elhage et al., 2022; Cunningham et al., 2023; Bricken et al., 2023).
Sparse Autoencoders Given an input $\bm{x}\in\mathbb{R}^{m}$, an SAE computes a sparse code using an encoder $\bm{z}=\sigma({\bm{W}}^{\top}(\bm{x}-\bm{b}_{\text{pre}})+\bm{b})$ and reconstructs the data as $\hat{\bm{x}}={\bm{D}}\bm{z}+\bm{b}_{\text{pre}}$, where ${\bm{W}}\in\mathbb{R}^{m\times p}$ and $\bm{b}\in\mathbb{R}^{p}$ are the encoder parameters, $\bm{b}_{\text{pre}}\in\mathbb{R}^{m}$ is a bias, and ${\bm{D}}\in\mathbb{R}^{m\times p}$ is the learned dictionary of interest with normalized atoms (i.e., $\|{\bm{D}}_{i}\|_{2}=1$). The nonlinearity $\sigma(\cdot)$ enforces sparsity; common choices include ReLU, TopK, and JumpReLU. These models are trained to minimize a sparsity-augmented reconstruction loss: $\mathcal{L}=\|\bm{x}-\hat{\bm{x}}\|_{2}^{2}+\lambda\,\mathcal{R}(\bm{z})+\alpha\,\mathcal{L}_{\text{aux}}$, where $\mathcal{R}(\bm{z})$ promotes sparsity, e.g., via $\ell_{1}$ regularization for ReLU (Cunningham et al., 2023; Bricken et al., 2023) or target-$\ell_{0}$ penalties (Rajamanoharan et al., 2024).
Orthogonality and Limitations of Shallow Recovery Sparse recovery (i.e., the inner problem optimization) in one iteration is theoretically guaranteed only when the dictionary ${\bm{D}}$ is sufficiently incoherent, i.e., when the mutual coherence $\mu({\bm{D}})=\max_{i\neq j}|{\bm{D}}_{i}^{\top}{\bm{D}}_{j}|$ is small (Makhzani & Frey, 2014; Arora et al., 2015). However, concepts underlying natural data may exhibit high coherence. This limits the ability of shallow, one-shot inference, sparse autoencoders to extract coherent concepts, and motivates the use of unrolled sparse-coding-based networks (Rambhatla et al., 2018; Tolooshams & Ba, 2022).
Figure 2: Expressivity.
3 Unrolling Matching Pursuit into MP-SAE
We propose Matching Pursuit Sparse Autoencoder (MP-SAE), an iterative inference procedure by unrolling Matching Pursuit (MP) (Mallat & Zhang, 1993) into a sparse autoencoder architecture. Unlike shallow SAEs, solving the inner problem in a single forward pass, MP-SAE performs inference sequentially.
Iterative Inference via Residual Updates  Matching Pursuit starts with a residual, defined as the difference between the input $\bm{x}$ and its current reconstruction $\hat{\bm{x}}$, i.e., $\bm{r}=\bm{x}-\hat{\bm{x}}$. Initially, this residual equals the input (or $\bm{x}-\bm{b}_{\text{pre}}$). MP-SAE iteratively reduces this residual by adding more atoms.
At each iteration, MP greedily selects the dictionary atom that best aligns with the current residual. This is done by computing the inner product between the residual and each atom, and selecting the one with the highest projection. Once the best-matching atom is selected, the algorithm projects the residual onto that atom to determine its contribution to the reconstruction. This contribution is then added to the current approximation of the input and subtracted from the residual. Over time, this greedy procedure iteratively reduces the residual and improves the reconstruction, step by step (see Figure 1 and Algorithm 1).
Theoretical Properties of Matching Pursuit  At each step, the residual $\bm{r}^{(t)}$ is orthogonal to the selected atom ${\bm{D}}_{j^{(t)}}$ (Proposition 8.1). Moreover, the norm of the residual decreases monotonically at each iteration (Proposition 8.2) and converges asymptotically to the component of the input orthogonal to the span of the dictionary ${\bm{D}}$ (Proposition 8.3).
Figure 3: Feature Selection vs. Activation Levels. Top: atoms with highest activation frequency ($\ell_{0}$). Bottom: atoms with highest activation $\mathbb{E}[{\bm{z}}_{j}]$ ($\ell_{1}$).
Unlike TopK that chooses all activations at once, MP infer sequentially: it selects one atom at a time, removes its contribution from the residual, and continues. Because each selected atom is subtracted from the residual, the algorithm naturally explores new directions —each selection is driven by what remains unexplained, pushing the model toward atoms that are less redundant and more complementary. This residual-driven mechanism improves diversity in the selected features and enhances robustness in the presence of dictionary coherence. Indeed, MP can achieve exponential convergence and accurately recover active components even when the dictionary is highly coherent, as long as it satisfies a block-incoherence condition, where correlations are localized within small groups of atoms (Peotta & Vandergheynst, 2007).
4 Results
We evaluate MP-SAE against four shallow SAEs: ReLU (Elhage et al., 2022; Cunningham et al., 2023), JumpReLU (Rajamanoharan et al., 2024), TopK (Gao et al., 2025), and BatchTopK (Bussmann et al., 2024), using the MNIST dataset. Additional results on large vision models, including expressivity and coherence, are provided in Appendix 7.
Expressivity We assess reconstruction performance by varying two key parameters: the sparsity level $k$ and the dictionary size $p$. Figure 2(a) shows that, when fixing $p=1000$ and varying $k$, MP-SAE is consistently more expressive across sparsity levels—despite having half the capacity of other SAEs due to the absence of encoder parameters from weight tying. As shown in Figure 2(b), when $k=10$ is fixed and $p$ is swept, MP-SAE continues to improve as the dictionary size grows, while shallow SAEs plateau. This highlights the efficiency of MP-SAE in leveraging additional capacity under the same sparsity constraint.
Learned Concepts In the following, we focus on dictionaries trained with $k=10$ and $p=1000$. The top row in Figure 3 shows 25 atoms with highest activation frequency for each SAE. All methods except ReLU appear to learn pen-stroke-like patterns, while ReLU learns atoms resembling full digits. The pen strokes learned by MP-SAE appear more precise. Interestingly, all shallow methods include a “negative” atom that closely resembles $-\bm{b}_{\text{pre}}$, which is the most frequently activated atom in ReLU and JumpReLU.
Figure 4: Activation distributions.
When sorting atoms by their average activation value $\mathbb{E}({\bm{z}}_{j})$ rather than frequency $\sum_{i}{\bm{z}}^{i}_{j}\neq 0$, a structural shift unique to MP-SAE emerges. As shown in the bottom row of Figure 3, the most heavily weighted atoms in MP-SAE resemble clean, idealized digit prototypes, in contrast to the more detailed pen strokes observed among its most frequently activated atoms. In comparison, shallow SAEs exhibit little variation between these two rankings; ReLU continues to activate full-digit atoms, while the others still primarily activate stroke-like patterns.
Hints of Hierarchy To capture this shift, the distributions of activation frequency and average activation value $\mathbb{E}[{\bm{z}}_{j}]$ are shown in Figure 4(a). JumpReLU, TopK, and BatchTopK exhibit high variance in activation frequency—some atoms are rarely used, while others are activated very frequently. In contrast, the variance in activation values remains low, with slightly higher values observed for the more frequently used atoms. By comparison, ReLU activates its dictionary atoms uniformly, both in frequency and magnitude. MP displays a perpendicular trend to the other $\ell_{0}$-based methods and JumpReLU: its atoms are activated with roughly equal frequency—similar to ReLU—but their activation values vary widely. Some atoms contribute much more to the reconstruction than others. A subtle inverse relationship is also observed: atoms with higher activation values tend to be used less frequently.
Figure 5: Coherence analysis of learned concepts.
This hierarchical behavior is supported by Figure 4(b), which shows that atoms with higher average activation values tend to be selected in the early layers of MP-SAE (the first iterations of Matching Pursuit). These atoms correspond to the digit-like patterns in the bottom row of Figure 3, and are refined by later atoms capturing more localized pen strokes (top row). This progression suggests a hierarchical structure in MP-SAE, building reconstructions from coarse to fine features (see Appendix 6).
To enable comparison with shallow encoders, atoms are reordered by activation value to simulate a sequential selection. $\ell_{0}$-based methods show a concentration of atoms toward the end, likely due to the auxiliary loss. This also highlights that ReLU exhibits an amplitude bias.
Coherence Finally, we assess coherence for both the learned dictionary and the atoms selected at inference. To move beyond pairwise similarity captured by mutual coherence, the Babel function (Tropp, 2004) is used (see Equation 2). It measures cumulative coherence $\mu_{1}(r)$, offering a more comprehensive view of redundancy within the dictionary. $\mu_{1}(r)$ reflects how well a single atom can be approximated by a group of $r$ others; lower values indicate lower redundancy.
Figure 5(a) shows that MP-SAE exhibits a more coherent dictionary than the shallow SAEs. However, as shown in Figure 5(b), it selects more incoherent atoms at inference. This highlights MP’s ability to draw incoherent subsets from a globally coherent dictionary. Interestingly, for shallow SAEs, the trends for the learned dictionary and the selected atoms align: ReLU consistently exhibits the highest coherence, while TopK remains the least coherent. This suggests that shallow SAEs are constrained to select more correlated atoms when the dictionary itself is more coherent.
5 Conclusion
We introduce MP-SAE and show through small-scale experiments on MNIST that it improves expressivity, learns hierarchical features, and overcomes coherence limitations of shallow SAEs. These experiments reveal distinct representation behaviors across SAEs and offer a foundation for building more interpretable sparse autoencoders.

### v2

@@
 This paper evaluates SAEs in a controlled setting using MNIST, which reveals that current shallow architectures implicitly rely on a quasi-orthogonality assumption that limits the ability to extract correlated features.
-To move beyond this, we introduce a multi-iteration SAE by unrolling Matching Pursuit (MP-SAE), enabling the residual-guided extraction of correlated features that arise in hierarchical settings such as handwritten digit generation while guaranteeing monotonic improvement of the reconstruction as more atoms are selected.
+To move beyond this, we compare them with an iterative SAE that unrolls Matching Pursuit (MP-SAE), enabling the residual-guided extraction of correlated features that arise in hierarchical settings such as handwritten digit generation while guaranteeing monotonic improvement of the reconstruction as more atoms are selected.
 Machine Learning, ICML
@@
 Sparse dictionary learning (Mairal et al., 2009; Rubinstein et al., 2010; Tosic & Frossard, 2011) aims to represent data ${\bm{x}}^{i}$ as sparse linear combinations of learned basis vectors ($\bm{D}$}, a.k.a atoms), for $i=1,\dots,n$, i.e.,
-${\bm{x}}^{i}\approx{\bm{D}}{\bm{z}}^{i}\quad\text{subject to}\quad\|{\bm{z}}^{i}\|_{0}\leq k,\quad\forall\ i=1,\dots,n.\vspace{-1.5mm}$ (1)
+${\bm{x}}^{i}\approx{\bm{D}}{\bm{z}}^{i}\quad\text{subject to}\quad\|{\bm{z}}^{i}\|_{0}\leq k,\quad\forall\ i=1,\dots,n.$ (1)
 where the constraint ensures that the sparse code ${\bm{z}}^{i}$ is at most $k$-sparse.
@@
 Building on this principle, recent work has proposed the use of sparse autoencoders (SAEs) to extract human-interpretable features from the internal activations of large language models (LLMs) (Elhage et al., 2022; Cunningham et al., 2023; Bricken et al., 2023; Rajamanoharan et al., 2024; Gao et al., 2025).
-These methods are motivated by the linear representation hypothesis  (Arora et al., 2016; Olah et al., 2020; Park et al., 2024) and the superposition hypothesis  (Elhage et al., 2022), i.e., internal representations can be modeled as sparse linear combinations of semantic directions.
+These methods are motivated by the linear representation hypothesis  (Arora et al., 2016; Park et al., 2024; Elhage et al., 2022), which posits that internal representations can be modeled as sparse linear combinations of semantic directions.
+Unlike the classical dictionary learning methods described above, SAEs solve the inner problem in a single iteration.
 Algorithm 1 Matching Pursuit Sparse Autoencoders (MP-SAE)
@@
 Among the few studies, (Hindupur et al., 2025) demonstrate that SAEs impose structural assumptions that shape what they can and cannot detect.
-Our Contributions In this work, we revisit the relationship between SAEs and dictionary learning by testing modern architectures on MNIST, a widely used benchmark in sparse coding (Aharon et al., 2006; Makhzani & Frey, 2014).
+Our Contributions In this work, we revisit the relationship between SAEs and dictionary learning by testing common interpretability SAEs architectures on MNIST, a controlled and well-established dataset in sparse coding literature.
 Despite similarities in architecture, we find that SAEs with different sparsity mechanisms yield structurally distinct dictionaries.
@@
 However, concepts underlying natural data may exhibit high coherence.
-This limits the ability of shallow, one-shot inference, sparse autoencoders to extract coherent concepts, and motivates the use of unrolled sparse-coding-based networks (Rambhatla et al., 2018; Tolooshams & Ba, 2022).
+This limitation motivates the use of unrolled sparse coding networks, which address the inner problem through multiple iterative updates, progressively refining the estimate of the sparse code $z$ as in classical sparse recovery methods.
+Shallow SAEs can be seen as approximating unrolled networks by performing only a single iteration: they rely on the same nonlinearities – JumpReLU implementing hard thresholding as in NOODL (Rambhatla et al., 2018), and ReLU implementing soft thresholding as in PUDLE (Tolooshams & Ba, 2022) – but do not perform iterative updates.
+3 Unrolling Matching Pursuit into MP-SAE
+We propose the Matching Pursuit Sparse Autoencoder (MP-SAE), which unrolls Matching Pursuit (MP) (Mallat & Zhang, 1993) into an iterative sparse autoencoder architecture.
+Previously used for image restoration tasks under the name Learned-MP (Khatib et al., 2021), our work focuses on the dictionary properties learned by MP-SAE and compares them to those obtained with shallow SAEs.
+Iterative Inference via Residual Updates  MP-SAE can be seen as an unrolled variant of TopK.
+Instead of selecting $k$ atoms in a single iteration, as in TopK, it selects one atom at each of the $k$ iterations.
+This allows the model to subtract the contribution of each selected atom from the input before the next selection, so that subsequent atoms are chosen to explain what remains unexplained.
+This unexplained part is captured by the residual, defined as the difference between the input $\bm{x}$ and its current reconstruction $\hat{\bm{x}}$, i.e., $\bm{r}=\bm{x}-\hat{\bm{x}}$.
+Initially, the residual is equal to the input (or $\bm{x}-\bm{b}_{\text{pre}}$).
 Figure 2: Expressivity.
-3 Unrolling Matching Pursuit into MP-SAE
-We propose Matching Pursuit Sparse Autoencoder (MP-SAE), an iterative inference procedure by unrolling Matching Pursuit (MP) (Mallat & Zhang, 1993) into a sparse autoencoder architecture.
-Unlike shallow SAEs, solving the inner problem in a single forward pass, MP-SAE performs inference sequentially.
-Iterative Inference via Residual Updates  Matching Pursuit starts with a residual, defined as the difference between the input $\bm{x}$ and its current reconstruction $\hat{\bm{x}}$, i.e., $\bm{r}=\bm{x}-\hat{\bm{x}}$.
-Initially, this residual equals the input (or $\bm{x}-\bm{b}_{\text{pre}}$).
-MP-SAE iteratively reduces this residual by adding more atoms.
 At each iteration, MP greedily selects the dictionary atom that best aligns with the current residual.
@@
 Moreover, the norm of the residual decreases monotonically at each iteration (Proposition 8.2) and converges asymptotically to the component of the input orthogonal to the span of the dictionary $\bm{D}$} (Proposition 8.3).
-Figure 3: Feature Selection vs. Activation Levels. Top: atoms with highest activation frequency ($\ell_{0}$). Bottom: atoms with highest activation $\mathbb{E}[{\bm{z}}_{j}]$ ($\ell_{1}$).
-Unlike TopK that chooses all activations at once, MP infer sequentially: it selects one atom at a time, removes its contribution from the residual, and continues.
-Because each selected atom is subtracted from the residual, the algorithm naturally explores new directions —each selection is driven by what remains unexplained, pushing the model toward atoms that are less redundant and more complementary.
+Because each selected atom is subtracted from the residual at each iteration, the algorithm naturally explores new directions – each selection is driven by what remains unexplained, pushing the model toward atoms that are less redundant and more complementary.
 This residual-driven mechanism improves diversity in the selected features and enhances robustness in the presence of dictionary coherence.
@@
 Learned Concepts In the following, we focus on dictionaries trained with $k=10$ and $p=1000$.
-The top row in Figure 3 shows 25 atoms with highest activation frequency for each SAE.
+The top row in Figure 3 shows 25 atoms with the highest activation frequency for each SAE.
 All methods except ReLU appear to learn pen-stroke-like patterns, while ReLU learns atoms resembling full digits.
 The pen strokes learned by MP-SAE appear more precise.
-Interestingly, all shallow methods include a “negative” atom that closely resembles $-\bm{b}_{\text{pre}}$, which is the most frequently activated atom in ReLU and JumpReLU.
-Figure 4: Activation distributions.
+Interestingly, all shallow methods include a “negative” atom that closely resembles $\bm{b}_{\text{pre}}$, which is the most frequently activated atom in ReLU and JumpReLU.
 When sorting atoms by their average activation value $\mathbb{E}({\bm{z}}_{j})$ rather than frequency $\sum_{i}{\bm{z}}^{i}_{j}\neq 0$, a structural shift unique to MP-SAE emerges.
@@
 In comparison, shallow SAEs exhibit little variation between these two rankings; ReLU continues to activate full-digit atoms, while the others still primarily activate stroke-like patterns.
+Figure 3: Feature Selection vs. Activation Levels. Top: atoms with highest activation frequency ($\ell_{0}$). Bottom: atoms with highest activation $\mathbb{E}[{\bm{z}}_{j}]$ ($\ell_{1}$).
 Hints of Hierarchy To capture this shift, the distributions of activation frequency and average activation value $\mathbb{E}[{\bm{z}}_{j}]$ are shown in Figure 4(a).
@@
 A subtle inverse relationship is also observed: atoms with higher activation values tend to be used less frequently.
-Figure 5: Coherence analysis of learned concepts.
 This hierarchical behavior is supported by Figure 4(b), which shows that atoms with higher average activation values tend to be selected in the early layers of MP-SAE (the first iterations of Matching Pursuit).
-These atoms correspond to the digit-like patterns in the bottom row of Figure 3, and are refined by later atoms capturing more localized pen strokes (top row).
-This progression suggests a hierarchical structure in MP-SAE, building reconstructions from coarse to fine features (see Appendix 6).
+These atoms correspond to digit-like patterns in the bottom row of Figure 3, and are refined by later atoms capturing more localized pen strokes (top row).
+This progression suggests a hierarchical structure in MP-SAE, reconstructing from coarse to fine features (see Appendix 6).
 To enable comparison with shallow encoders, atoms are reordered by activation value to simulate a sequential selection.
@@
 This also highlights that ReLU exhibits an amplitude bias.
-Coherence Finally, we assess coherence for both the learned dictionary and the atoms selected at inference.
-To move beyond pairwise similarity captured by mutual coherence, the Babel function (Tropp, 2004) is used (see Equation 2).
-It measures cumulative coherence $\mu_{1}(r)$, offering a more comprehensive view of redundancy within the dictionary.
-$\mu_{1}(r)$ reflects how well a single atom can be approximated by a group of $r$ others; lower values indicate lower redundancy.
+Figure 4: Activation distributions.
+Coherence Finally, we assess the coherence of both the learned dictionary and the atoms selected at inference time using the Babel score (Tropp, 2004).
+Unlike mutual coherence, which captures only the maximum pairwise similarity between atoms, the Babel function provides a more comprehensive measure of redundancy by quantifying cumulative coherence.
+Given a dictionary ${\bm{D}}=[{\bm{D}}_{1},\dots,{\bm{D}}_{p}]\in\mathbb{R}^{m\times p}$ with unit-norm columns, the Babel function of order $r$ is defined as:
+$\mu_{1}(r)=\max_{\begin{subarray}{c}S\subset[p],|S|=r\end{subarray}}\left(\max_{j\notin S}\sum_{i\in S}\left|{\bm{D}}_{i}^{\top}{\bm{D}}_{j}\right|\right)$
+The value $\mu_{1}(r)$ reflects how well a single atom can be approximated by a group of $r$ others—lower values indicate lower redundancy. Further details are provided in Appendix 7.
+Figure 5: Coherence analysis of learned concepts.
 Figure 5(a) shows that MP-SAE exhibits a more coherent dictionary than the shallow SAEs.
@@
 We introduce MP-SAE and show through small-scale experiments on MNIST that it improves expressivity, learns hierarchical features, and overcomes coherence limitations of shallow SAEs.
-These experiments reveal distinct representation behaviors across SAEs and offer a foundation for building more interpretable sparse autoencoders.
+These experiments also reveal distinct representation behaviors across SAEs and offer a foundation for building more interpretable sparse autoencoders.